OpenStudy (anonymous):

write an equation in point-slope form for each: a line with a slope of 4 that passes through point (3,6). Show work. plz

OpenStudy (anonymous):

Point slope form is y-y1=m(x-x1) Now all you have to do is plugin your variables so your final equation would be: y-6=4(x-3) Hope that helps!

OpenStudy (anonymous):

4x-y=6

OpenStudy (anonymous):

y-6=4(x-3) is already in point slope form. You dont have to do anything else to it